# Attractors for the Wave Equation with Nonlinear Energy Critical Sources and Dissipation

### v2

@@ 4.3 Gradient System
Assume that the assumptions of Theorem 2.2 hold. Then, $\varphi:\mathcal{H}\to\mathbb{R}$ given by
[…]
is a strict Lyapunov functional for the dynamical system $(\mathcal{H},S_{t})$. Consequently, $(\mathcal{H},S_{t})$ is a gradient dynamical system.
@@
Let us define $\varphi:=E$. From energy equality (2.8), we have
$\varphi(S_{t}U_{0})+\int_{0}^{t}\int_{\Omega}g(u_{t})u_{t}dxd\tau=\varphi(U_{0}),$ (4.21)
for every $U_{0}\in\mathcal{H}$. From Assumption (2.1), we have
$\int_{0}^{t}\int_{\Omega}g(u_{t})u_{t}dxd\tau\geq\kappa_{0}\int_{0}^{t}\int_{\Omega}|u_{t}|^{6}dxd\tau.$ (4.22)
Then, it follows from (4.21) and (4.22) that $\varphi(S_{t}U_{0})=\varphi(U_{0})$. Hence
$\int_{0}^{t}\|u_{t}(\tau)\|^{6}_{6}\,d\tau\leq 0.$
Using embedding $L^{6}(\Omega)\hookrightarrow H^{0}$, we conclude that
@@
then $\varphi$ is a strict Lyapunov functional for the dynamical system $(\mathcal{H},S_{t})$. ∎
5 Proof of Theorem 2.5: Quasi-stability property
The existence of a compact global attractor $\mathfrak{A}$ is guaranteed by Theorem 2.4 under weaker assumption $g^{\prime}(0)\geq 0$. However, a strict inequality $g^{\prime}(0)>0$ allows to show that the dynamical system $(\mathcal{H},S_{t})$ is quasi-stable according to Definition 7.8. And within the context of the theory of dynamical systems associated with the study of attractors, for quasi-stable systems, it is possible to obtain important properties such as the finite dimension and smoothness of the attractor, as well as the existence of exponential attractors. Properties that will be established for the dynamical system associated with problem (LABEL:P) in Theorem 2.6.
The proof that the system $(\mathcal{H},S_{t})$ is quasi-stable follows from the fact that $(\mathcal{H},S_{t})$ satisfies conditions (7.1) and (7.2) of the Definition 7.8 of a quasi-stable system. The condition (7.1) is guaranteed directly from the Lipschitz property (3.52). It remains to prove the quasi-stability inequality (7.2) which is ensured by the relation (5.1) established in the following proposition.
5.1 Quasi-stability Estimate
@@
$g^{\prime}(s)\geq\kappa_{2},\quad\mbox{for all}\quad s\in\mathbb{R}.$ (5.2)
Therefore, from (5.2) and using the same arguments used in [18, Remark 4:1] with $\gamma=4$, we obtain
$\left[\,g(r)-g(s)\,\right](r-s)\geq\kappa_{2}|s-r|^{2}+\frac{\kappa_{0}}{10}\left(|r|^{4}+|s|^{4}\right)|r-s|^{2}.$ (5.3)
@@
$\left\{\begin{array}[]{l}\displaystyle{z_{tt}-\Delta z+\left[\,g(u^{1}_{t})-g(u^{2}_{t})\,\right]+\left[\,f(u^{1})-f(u^{2})\,\right]=0,\quad\mbox{in}\quad\Omega\times[0,\infty),}\\ \displaystyle{(z(0),z_{t}(0))=(u_{0}-v_{0},u_{1}-v_{1}),\quad z=0,\quad\mbox{on}\quad\Gamma.}\end{array}\right$. (5.6)
Using the energy identity (2.8) for the difference of two weak solutions, we have
$\begin{aligned} \frac{d}{dt}E_{z}(t)+\int_{\Omega}D(z_{t})dx=&\;K_{f}\int_{\Omega}zz_{t}dx\\ &+\frac{1}{2}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{1}+(1-\theta)u^{2})\left[\theta u^{1}_{t}+(1-\theta)u^{2}_{t}\right]d\theta|z|^{2}dx.\end{aligned}$ (5.7)
where
$D(z_{t}):=\left[g(u^{1}_{t})-g(u^{2}_{t})\right]z_{t}$
and
@@
Integrating (5.7) from $\tau\geq 0$ to $T$, we get
$\begin{aligned} &E_{z}(T)-E_{z}(\tau)+\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ &=\underbrace{K_{f}\int_{\tau}^{T}\int_{\Omega}zz_{t}ds}_{J_{1}}+\underbrace{\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\theta u^{1}+(1-\theta)u^{2})\left[\theta u^{1}_{t}+(1-\theta)u^{2}_{t}\right]d\theta|z|^{2}dxds}_{J_{2}}.\end{aligned}$ (5.8)
@@
$|J_{1}|$ $\leq$ $K_{f}\int_{\tau}^{T}\|z\|\|z_{t}\|ds\leq\frac{K_{f}}{\kappa_{2}^{1/2}}\left(\int_{\tau}^{T}\|z\|^{2}ds\right)^{1/2}\left(\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds\right)^{1/2}$
$\leq$ $\frac{K_{f}^{2}}{2\kappa_{2}}\int_{\tau}^{T}\|z\|^{2}ds+\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds.$
@@
$|J_{2}|$ $\leq$ $\;C\int_{\tau}^{T}\left[\,1+\|u^{1}\|^{3}_{6}+\|u^{2}\|^{3}_{6}\,\right]\left[\,\|u^{1}_{t}\|_{6}+\|u^{2}_{t}\|_{6}\,\right]\|z\|_{6}^{2}ds$
$\leq$ $C_{B}\int_{\tau}^{T}\left[\,\|u^{1}_{t}\|_{6}+\|u^{2}_{t}\|_{6}\,\right]\|\nabla z\|^{2}ds$
$\leq$ $\delta\int_{\tau}^{T}E_{z}(s)ds+C_{B,\delta}\int_{\tau}^{T}d(s,u^{1},u^{2})E_{z}(s)ds,$
where
$d(t,u^{1},u^{2}):=\|u^{1}_{t}(t)\|^{6}_{6}+\|u^{2}_{t}(t)\|^{6}_{6}.$
Substituting $|J_{1}|$ and $|J_{2}|$ into (LABEL:ASP-11), we find
$\begin{aligned} E_{z}(T)-E_{z}(\tau)+\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds\leq&\;\frac{K_{f}^{2}}{2\kappa_{2}}\int_{\tau}^{T}\|z\|^{2}ds+\delta\int_{\tau}^{T}E_{z}(s)ds\\ &+C_{B,\delta}\int_{\tau}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.9)
@@
$\begin{aligned} TE_{z}(T)+\frac{1}{2}\int_{0}^{T}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxdsd\tau\leq&\;\int_{0}^{T}E(\tau)d\tau+\frac{K_{f}^{2}}{2\kappa_{2}}\int_{0}^{T}\int_{\tau}^{T}\|z\|^{2}dsd\tau\\ &+\delta T\int_{0}^{T}E_{z}(s)ds+TC_{B,\delta}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.10)
@@
$\begin{aligned} 2\int_{0}^{T}E_{z}(s)ds=&\;2\int_{0}^{T}\|z_{t}\|^{2}ds+K_{f}\int_{0}^{T}\|z\|^{2}ds\\ &\;-\underbrace{\left[\int_{\Omega}z_{t}(T)z(T)dx-\int_{\Omega}z_{t}(0)z(0)dx\right]}_{J_{3}}-\underbrace{\int_{0}^{T}\int_{\Omega}\left[g(u^{1}_{t})-g(u^{2}_{t})\right]zdxds}_{J_{4}}.\end{aligned}$ (5.11)
@@
$2\int_{0}^{T}\|z_{t}\|^{2}ds\leq\frac{2}{\kappa_{2}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds.$ (5.12)
@@
$|-J_{3}|\leq\frac{1}{\lambda_{1}^{1/2}}\left[\,E_{z}(T)+E_{z}(0)\,\right].$ (5.13)
Still, using equality (LABEL:ASP-11) and estimates $|\mathrm{J}_{1}|$ and $|\mathrm{J}_{2}|$ (with $\tau=0$) and choosing $\delta=\frac{\lambda_{1}^{1/2}}{2}$, we obtain
$\begin{aligned} E_{z}(0)\leq&\;E_{z}(T)+\frac{3}{2}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{K^{2}_{f}}{2\kappa_{2}}\int_{0}^{T}\|z\|^{2}ds\\ &+\frac{\lambda_{1}^{1/2}}{2}\int_{0}^{T}E_{z}(s)ds+C_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.14)
@@
$\begin{aligned} |-J_{3}|\leq&\;\frac{2}{\lambda_{1}^{1/2}}E_{z}(T)+\frac{3}{2\lambda_{1}^{1/2}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{K^{2}_{f}}{2\kappa_{2}\lambda_{1}^{1/2}}\int_{0}^{T}\|z\|^{2}ds\\ &+\frac{1}{2}\int_{0}^{T}E_{z}(s)ds+C_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.15)
@@
$\begin{aligned} |-J_{4}|=&\;\left|-\int_{0}^{T}\int_{\Omega}\int_{0}^{1}g^{\prime}(\theta u^{1}_{t}+(1-\theta)u^{2}_{t})d\theta z_{t}zdxds\right|\\ \leq&\;\kappa_{1}\int_{0}^{T}\int_{\Omega}\int_{0}^{1}\left[\,1+|\theta u^{1}_{t}+(1-\theta)u^{2}_{t}|^{4}\,\right]d\theta|z_{t}||z|dxds\\ \leq&\;\kappa_{1}\int_{0}^{T}\|z_{t}\|\|z\|ds+8\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}||z|dxds\\ \leq&\;\int_{0}^{T}\|z_{t}\|^{2}ds+\frac{\kappa_{1}^{2}}{4}\int_{0}^{T}\|z\|^{2}ds+8\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}||z|dxds\\ \leq&\;\frac{1}{\kappa_{2}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{\kappa_{1}^{2}}{4}\int_{0}^{T}\|z\|^{2}ds+\underbrace{8\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}||z|dxds}_{J_{4}^{\prime}}\end{aligned}$ (5.16)
@@
$\begin{aligned} J^{\prime}_{4}\leq&\;8\kappa_{1}\int_{0}^{T}\left(\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dx\right)^{1/2}\left(\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z|^{2}dx\right)^{1/2}ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\left[\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dx+\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z|^{2}dx\right]\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds\\ &+4\kappa_{1}\int_{0}^{T}\left(\int_{\Omega}\left[\,|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\,\right]^{3/2}dx\right)^{2/3}\left(\int_{\Omega}|z|^{6}dx\right)^{1/3}ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds+4\kappa_{1}\int_{0}^{T}\left[\,|u^{1}_{t}\|^{4}_{6}+\|u^{2}_{t}\|^{4}_{6}\,\right]\|z\|^{2}_{6}ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds+\frac{4\kappa_{1}}{\lambda}\int_{0}^{T}\left[\,|u^{1}_{t}\|^{4}_{6}+\|u^{2}_{t}\|^{4}_{6}\,\right]\|\nabla z\|^{2}ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds+\frac{8\kappa_{1}}{\lambda}\int_{0}^{T}\left[\,|u^{1}_{t}\|^{4}_{6}+\|u^{2}_{t}\|^{4}_{6}\,\right]E_{z}(s)ds\\ \leq&\;4\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}|^{2}dxds+\frac{32\kappa_{1}^{2}}{\lambda^{2}}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+\frac{1}{2}\int_{0}^{T}E_{z}(s)ds\\ \leq&\;\frac{40\kappa_{1}}{\kappa_{0}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{32\kappa_{1}^{2}}{\lambda^{2}}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+\frac{1}{2}\int_{0}^{T}E_{z}(s)ds.\end{aligned}$ (5.17)
@@
$\begin{aligned} \int_{0}^{T}E_{z}(s)ds\leq&\;\frac{2}{\lambda_{1}^{1/2}}E_{z}(T)+C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ &+C_{1}\int_{0}^{T}\|z\|^{2}ds+C^{\prime}_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds,\end{aligned}$ (5.18)
where $C_{0}=\frac{2}{\kappa_{2}}+\frac{3}{2\lambda_{1}^{1/2}}+\frac{1}{\kappa_{2}}+\frac{40\kappa_{1}}{\kappa_{2}}$, $C_{1}=K_{f}+\frac{K_{f}^{2}}{2\kappa\lambda_{1}^{1/2}}+\frac{\kappa_{1}^{2}}{4}$, and $C^{\prime}_{B}=C_{B}+\frac{32\kappa_{1}^{2}}{\lambda^{2}}$. Combining (5.10) and (5.18) we obtain that
$\begin{aligned} \left(T-\frac{4}{\lambda_{1}^{1/2}}\right)E_{z}(t)+\left(1-\delta T\right)\int_{0}^{T}E(s)ds\leq\;2C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ +\left(TC_{B,\delta}+2C^{\prime}_{B}\right)\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)\int_{0}^{T}\|z\|^{2}ds.\end{aligned}$ (5.19)
Using without loss of generality that $\frac{T}{2}-\frac{4}{\lambda_{1}^{1/2}}>0$ and taking $\delta$ small enough such that $1-\delta T\geq\frac{1}{2}$, it follows from (5.19) that
$\begin{aligned} TE_{z}(T)+\int_{0}^{T}E(s)ds\leq\;4C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ +2\left(TC_{B,\delta}+2C^{\prime}_{B}\right)\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+2\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)\int_{0}^{T}\|z\|^{2}ds.\end{aligned}$ (5.20)
@@
$\begin{aligned} \frac{1}{2}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\leq&\;E_{z}(0)-E_{z}(T)+\delta\int_{0}^{T}E_{z}(s)ds\\ &+C_{B,\delta}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$ (5.21)
Substituting (5.21) into (5.20) and choosing $\delta=\frac{1}{2C_{0}}$, we obtain
$E_{z}(T)$ $\leq$ $\frac{2C_{0}}{T+2C_{0}}E_{z}(0)+\frac{4C_{0}\left(TC_{B,\delta}+2C^{\prime}_{B}\right)}{T+2C_{0}}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds$
$+\,\frac{4C_{0}\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)}{T+2C_{0}}\sup_{s\in[0,T]}\|z(s)\|^{2}.$
@@
with $0<\gamma\equiv\frac{2C_{0}}{T+2C_{0}}<1$ and $C_{B}=\frac{4C_{0}\left(TC_{B,\delta}+2C^{\prime}_{B}\right)}{T+2C_{0}}+\frac{4C_{0}\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)}{T+2C_{0}}$, where
$b_{m}\equiv\sup_{s\in[mT,(m+1)T]}\|z(s)\|^{2}+\int_{mT}^{(m+1)T}d(s;u^{1},u^{2})E_{z}(s)ds.$

### v4

@@ 4.2. Gradient System
Assume that the assumptions of Theorem 2.2 hold. Then, $(\mathcal{H},S_{t})$ is a gradient dynamical system.
@@
Let us take $\Phi$ as the energy functional $E$ defined in (2.9). From energy equality (2.8), we have
(4.6) $\Phi(S_{t}U_{0})+\int_{0}^{t}\int_{\Omega}g(u_{t})u_{t}dxd\tau=\Phi(U_{0}),$
for every $U_{0}\in\mathcal{H}$. Hence $\Phi(S_{t}U_{0})$ is non increasing. Now let us suppose $\Phi(S_{t}U_{0})=\Phi(U_{0})$ for all $t\geq 0$. Then from above From Assumption (2.1), we have
(4.7) $0\leq\kappa_{0}\int_{0}^{t}\int_{\Omega}|u_{t}|^{6}dxd\tau\leq\int_{0}^{t}\int_{\Omega}g(u_{t})u_{t}dxd\tau=0.$
Then, using embedding $L^{6}(\Omega)\hookrightarrow H^{0}$, it follows from (4.7) that
@@
then $\Phi$ is a strict Lyapunov functional for the dynamical system $(\mathcal{H},S_{t})$. ∎
4.3. Boundedness of the set of stationary solutions
Proposition 4.4.
Let $\Phi$ be the Lyapunov function given in Proposition 4.3. Then $\Phi(U)$ is bounded from above on any bounded subset of $\mathcal{H}$ and the set $\Phi_{R}=\{U:\Phi(U)\leq R\}$ is bounded for every $R$.
Proof.
Let $R>0$ and we define $\Phi_{R}=\{U\in\mathcal{H}|\Phi(U)\leq R\}.$ From Assumption (2.5) and Hölder’s inequality, it follows from (4.6) that
[…]
Therefore, for $U(t)\in\Phi_{R}$, we get
[…]
Which implies that $\Phi_{R}$ is bounded for every $R>0$. ∎
Proposition 4.5.
The set $\mathcal{N}$ of stationary solutions of problem (LABEL:P) is bounded in $\mathcal{H}.$
Proof.
(4.8) $-\Delta u+f(u)=h,\quad u\in H^{1},$
(4.9) $\|\nabla u\|^{2}+\int_{\Omega}f(u)udx=\int_{\Omega}hudx.$
From Assumption (2.5) and Hölder inequality, it follows from (4.9) that
(4.10) $\|\nabla u\|^{2}\leq\frac{1}{\omega^{2}\lambda_{1}}\|h\|^{2}+\frac{2C_{\nu}|\Omega|}{\omega}.$
5. Proof of Theorem 2.5: Quasi-stability property
The existence of a compact global attractor $\mathfrak{A}$ is guaranteed by Theorem 2.4 under weaker assumption $g^{\prime}(0)\geq 0$. However, a strict inequality $g^{\prime}(0)>0$ allows to show that the dynamical system $(\mathcal{H},S_{t})$ is quasi-stable as defined in [16, Definition 7.9.2]. And within the context of the theory of dynamical systems associated with the study of attractors, for quasi-stable systems, it is possible to obtain important properties such as the finite dimension and smoothness of the attractor, as well as the existence of exponential attractors. Properties that will be established for the dynamical system associated with problem (LABEL:P) in Theorem 2.6.
The proof that the system $(\mathcal{H},S_{t})$ is quasi-stable follows from the fact that $(\mathcal{H},S_{t})$ satisfies conditions (7.9.2) and (7.9.3) from [16, Definition 7.9.2]. The condition (7.9.2) is guaranteed directly from the Lipschitz property (3.44). It remains to prove the quasi-stability inequality (7.9.3) which is ensured by the relation (5.1) established in the following proposition.
5.1. Quasi-stability Estimate
@@
(5.2) $g^{\prime}(s)\geq\kappa_{2},\quad\mbox{for all}\quad s\in\mathbb{R}.$
Therefore, from (5.2) and using the same arguments used in [20, Remark 4:1] with $\gamma=4$, we obtain
(5.3) $\left[\,g(r)-g(s)\,\right](r-s)\geq\kappa_{2}|s-r|^{2}+\frac{\kappa_{0}}{10}\left(|r|^{4}+|s|^{4}\right)|r-s|^{2}.$
@@
(5.6) $\left\{\begin{array}[]{l}\displaystyle{z_{tt}-\Delta z+\left[\,g(u^{1}_{t})-g(u^{2}_{t})\,\right]+\left[\,f(u^{1})-f(u^{2})\,\right]=0,\quad\mbox{in}\quad\Omega\times[0,\infty),}\\ \displaystyle{(z(0),z_{t}(0))=(u_{0}-v_{0},u_{1}-v_{1}),\quad z=0,\quad\mbox{on}\quad\Gamma.}\end{array}\right$.
Let $\chi_{\theta}(t):=\theta u^{1}(t)+(1-\theta)u^{2}(t)$. Using the energy identity (2.8) for the difference of two weak solutions and integrating the relation obtained from $\tau\geq 0$ to $T$, we get
(5.7) $E_{z}(T)-E_{z}(\tau)+\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds=\;\underbrace{K_{f}\int_{\tau}^{T}\int_{\Omega}zz_{t}ds}_{J_{1}}+\underbrace{\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}\int_{0}^{1}f^{\prime\prime}(\chi_{\theta})(\chi_{\theta})_{t}d\theta|z|^{2}dxds}_{J_{2}},$
where $D(z_{t}):=\left[g(u^{1}_{t})-g(u^{2}_{t})\right]z_{t}$ and
@@
$|J_{1}|\leq\frac{K_{f}^{2}}{2\kappa_{2}}\int_{\tau}^{T}\|z\|^{2}ds+\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds.$
@@
$|J_{2}|$ $\leq$ $\delta\int_{\tau}^{T}E_{z}(s)ds+C_{B,\delta}\int_{\tau}^{T}d(s,u^{1},u^{2})E_{z}(s)ds,$
where $d(t,u^{1},u^{2}):=\|u^{1}_{t}(t)\|^{6}_{6}+\|u^{2}_{t}(t)\|^{6}_{6}.$ Then, substituting $|J_{1}|$ and $|J_{2}|$ into (5.7), we find
(5.8) $\begin{aligned} E_{z}(T)-E_{z}(\tau)+\frac{1}{2}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxds\leq&\;\frac{K_{f}^{2}}{2\kappa_{2}}\int_{\tau}^{T}\|z\|^{2}ds+\delta\int_{\tau}^{T}E_{z}(s)ds\\ &+C_{B,\delta}\int_{\tau}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$
@@
(5.9) $\begin{aligned} &TE_{z}(T)+\frac{1}{2}\int_{0}^{T}\int_{\tau}^{T}\int_{\Omega}D(z_{t}(s))dxdsd\tau\;\leq\;\int_{0}^{T}E(\tau)d\tau\\ &+\frac{K_{f}^{2}}{2\kappa_{2}}\int_{0}^{T}\int_{\tau}^{T}\|z\|^{2}dsd\tau+\delta T\int_{0}^{T}E_{z}(s)ds+TC_{B,\delta}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$
@@
(5.10) $\begin{aligned} 2\int_{0}^{T}E_{z}(s)ds=&\;2\int_{0}^{T}\|z_{t}\|^{2}ds+K_{f}\int_{0}^{T}\|z\|^{2}ds\\ &-\underbrace{\left[\int_{\Omega}z_{t}(s)z(s)dx\right]^{s=T}_{s=0}}_{J_{3}}-\underbrace{\int_{0}^{T}\int_{\Omega}\left[g(u^{1}_{t})-g(u^{2}_{t})\right]zdxds}_{J_{4}}.\end{aligned}$
@@
(5.11) $2\int_{0}^{T}\|z_{t}\|^{2}ds\leq\frac{2}{\kappa_{2}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds.$
@@
(5.12) $|-J_{3}|\leq\frac{1}{\lambda_{1}^{1/2}}\left[\,E_{z}(T)+E_{z}(0)\,\right].$
Still, using equality (5.7) and estimates $|J_{1}|$ and $|J_{2}|$ (with $\tau=0$) and choosing $\delta=\frac{\lambda_{1}^{1/2}}{2}$, we obtain
(5.13) $\begin{aligned} E_{z}(0)\leq&\;E_{z}(T)+\frac{3}{2}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{K^{2}_{f}}{2\kappa_{2}}\int_{0}^{T}\|z\|^{2}ds\\ &+\frac{\lambda_{1}^{1/2}}{2}\int_{0}^{T}E_{z}(s)ds+C_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$
@@
(5.14) $\begin{aligned} |-J_{3}|\leq&\;\frac{2}{\lambda_{1}^{1/2}}E_{z}(T)+\frac{3}{2\lambda_{1}^{1/2}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{K^{2}_{f}}{2\kappa_{2}\lambda_{1}^{1/2}}\int_{0}^{T}\|z\|^{2}ds\\ &+\frac{1}{2}\int_{0}^{T}E_{z}(s)ds+C_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.\end{aligned}$
@@
(5.15) $|-J_{4}|\leq\;\frac{1}{\kappa_{2}}\int_{0}^{T}\int_{\Omega}D(z_{t})dxds+\frac{\kappa_{1}^{2}}{4}\int_{0}^{T}\|z\|^{2}ds+\underbrace{8\kappa_{1}\int_{0}^{T}\int_{\Omega}\left[|u^{1}_{t}|^{4}+|u^{2}_{t}|^{4}\right]|z_{t}||z|dxds}_{J_{4}^{\prime}}$
@@
(5.16) $J^{\prime}_{4}\leq$ $\;\frac{40\kappa_{1}}{\kappa_{0}}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds+\frac{32\kappa_{1}^{2}}{\lambda^{2}}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+\frac{1}{2}\int_{0}^{T}E_{z}(s)ds.$
@@
(5.17) $\begin{aligned} \int_{0}^{T}E_{z}(s)ds\leq&\;\frac{2}{\lambda_{1}^{1/2}}E_{z}(T)+C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ &+C_{1}\int_{0}^{T}\|z\|^{2}ds+C^{\prime}_{B}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds,\end{aligned}$
where $C_{0}=\frac{2}{\kappa_{2}}+\frac{3}{2\lambda_{1}^{1/2}}+\frac{1}{\kappa_{2}}+\frac{40\kappa_{1}}{\kappa_{2}}$, $C_{1}=K_{f}+\frac{K_{f}^{2}}{2\kappa\lambda_{1}^{1/2}}+\frac{\kappa_{1}^{2}}{4}$, and $C^{\prime}_{B}=C_{B}+\frac{32\kappa_{1}^{2}}{\lambda^{2}}$. Combining (LABEL:ASP-1111) and (5.17) we obtain that
(5.18) $\begin{aligned} &\left(T-\frac{4}{\lambda_{1}^{1/2}}\right)E_{z}(t)+\left(1-\delta T\right)\int_{0}^{T}E(s)ds\leq\;2C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ &\quad+\left(TC_{B,\delta}+2C^{\prime}_{B}\right)\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)\int_{0}^{T}\|z\|^{2}ds.\end{aligned}$
Using without loss of generality that $\frac{T}{2}-\frac{4}{\lambda_{1}^{1/2}}>0$ and taking $\delta$ small enough such that $1-\delta T\geq\frac{1}{2}$, it follows from (LABEL:ASP-33) that
(5.19) $\begin{aligned} &TE_{z}(T)+\int_{0}^{T}E(s)ds\leq\;4C_{0}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\\ &\quad+2\left(TC_{B,\delta}+2C^{\prime}_{B}\right)\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds+2\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)\int_{0}^{T}\|z\|^{2}ds.\end{aligned}$
@@
(5.20) $\frac{1}{2}\int_{0}^{T}\int_{\Omega}D(z_{t}(s))dxds\leq$ $\;E_{z}(0)-E_{z}(T)+\delta\int_{0}^{T}E_{z}(s)ds+C_{B,\delta}\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds.$
Substituting (5.20) into (LABEL:ASP-333) and choosing $\delta=\frac{1}{2C_{0}}$, we obtain
$E_{z}(T)\leq\frac{2C_{0}}{T+2C_{0}}E_{z}(0)+C_{B}\left[\sup_{s\in[0,T]}\|z(s)\|^{2}+\int_{0}^{T}d(s,u^{1},u^{2})E_{z}(s)ds\right],$
where $C_{B}=\frac{4C_{0}\left(TC_{B,\delta}+2C^{\prime}_{B}\right)}{T+2C_{0}}+\frac{4C_{0}\left(2C_{1}+\frac{K_{f}T}{2\kappa_{2}}\right)}{T+2C_{0}}$.
@@
where
$0<\gamma\equiv\frac{2C_{0}}{T+2C_{0}}<1\quad\mbox{and}\quad b_{m}\equiv\sup_{s\in[mT,(m+1)T]}\|z(s)\|^{2}+\int_{mT}^{(m+1)T}d(s;u^{1},u^{2})E_{z}(s)ds.$